I really need your help in solving this problem. I appreciate your time and efforts. Have a terrific day tomorrow :) Here is the question? Solve each equation:
6x^2 + 1=5x

Answer by rothauserc(4718)

6x^2 + 1 = 5x
subtract 5x from both sides of =
6x^2 -5x +1 = 0
factor the polynomial
(3x-1)*(2x-1) = 0
x = 1/3 and x = 1/2